The wave length and a set of wave travelling in a medium are 60cm and 320cm² means a plane refraction surface at an angle of 60°.If velocity after is 280cm², calculate the wave length of the wave in the second medium and also the angle of refraction.


1
Expert's answer
2021-01-20T08:43:13-0500

First, find the refractive index:


"n=\frac{v_i}{v_r}=\frac{320}{280}=1.14."

Next, find the angle of refraction using Snell's law:


"n=\frac{\text{ sin}\theta_i}{\text{ sin}\theta_r},\\\space\\
\theta_r=\text{arcsin}\frac{\text{ sin}\theta_i}{n}=\text{arcsin}\frac{\text{ sin}60°}{1.14}=49.4°."

Finally, find the wavelength of the refracted wave:


"\lambda_r=\frac{\lambda_i}{n}=\frac{60}{1.14}=52.6\text{ cm}."